OpenStudy (anonymous):

How can you write the expression with rationalized denominator? √3-√6/√3+√6

OpenStudy (campbell_st):

you need to multiply it by 1 written in an equivalent form use the same values as the denominator with the opposite operator between the terms \[\frac{\sqrt{3} - \sqrt{6}}{\sqrt{3} + \sqrt{6}} \times \frac{\sqrt{3}-\sqrt{6}}{\sqrt{3}-\sqrt{6}}\] completing the multiplication will give a rational number in the denominator and the numerator will be the expansion of a perfect square. hope it helps

OpenStudy (anonymous):

Okay, thank you!!